The ratio of girls to boys of a dance group is 5:3. Thee would had been equal numbr of boys and girls if thee had been 3 more boys and 3 fewer girls. How many girls and boys are in a dance group?
Answer by ankor@dixie-net.com(22740)

The ratio of girls to boys of a dance group is 5:3.
There would had been equal number of boys and girls if there had been 3 more boys and 3 fewer girls.
How many girls and boys are in a dance group?
:
Let x = the multiplier
then
5x = no. of girls
and
3x = no. or boys
"There would had been equal number of boys and girls if there had been 3 more boys and 3 fewer girls.
Girls = Boys
5x - 3 = 3x + 3
5x - 3x = 3 + 3
2x = 6
x = 3 is the multiplier
therefore
3*5 = 15 girls
and
3 * 3 = 9 boys